# Permanents of matrix ensembles: computation, distribution, and geometry

### v1

@@ Abstract.
We report on a computational and experimental study of permanents. On the computational side, we use the GPU to greaatly accelerate the computation of permanents over $\mathbb{C},$ $\mathbb{R},$ $\mathbb{F}_{p}$ and $\mathbb{Q}.$ In particular, we use this to compute the permanents of DFT and Schur matrices far beyond the ranges hitherto known. On the experimental side, we present two new observations. First, for Haar-distributed unitary matrices $U$, the permanent $\operatorname{perm}(U)$ follows a circularly-symmetric complex Gaussian distribution $\mathcal{CN}(0,\sigma^{2})$ — we confirm this via a number of tests for $n$ up to 23 with 50,00050{,}000 samples. The DFT matrix permanent is an extreme outlier for every prime $n\geq 7$. In contrast, for Haar-random orthogonal matrices $O$, the permanent $\operatorname{perm}(O)$ is approximately real Gaussian but with positive excess kurtosis that decays as $O(1/n)$, indicating slower convergence. For matrices with Gaussian entries (GUE, GOE, Ginibre), the permanent follows an $\alpha$-stable distribution with stability index $\alpha\approx 1.0$$1.4$, well below the Gaussian value $\alpha=2$.
@@
for odd $n$ and zero for even $n$. For the geodesic to the DFT matrix, the permanent recovers 10104040 times above its valley minimum when $n$ is prime, but not when $n$ is composite — a geodesic fingerprint of primality.

### v2

@@ Abstract.
We report on a computational and experimental study of permanents. On the computational side, we use the GPU to greaatly accelerate the computation of permanents over $\mathbb{C},$ $\mathbb{R},$ $\mathbb{F}_{p}$ and $\mathbb{Q}.$ First, for Haar-distributed unitary matrices $U$, the permanent $\operatorname{perm}(U)$ follows a circularly-symmetric complex Gaussian distribution $\mathcal{CN}(0,\sigma^{2})$ — we confirm this via a number of tests for $n$ up to 23 with 50,00050{,}000 samples. The DFT matrix permanent is an extreme outlier for every prime $n\geq 7$. In contrast, for Haar-random orthogonal matrices $O$, the permanent $\operatorname{perm}(O)$ is approximately real Gaussian but with positive excess kurtosis that decays as $O(1/n)$, indicating slower convergence. For matrices with Gaussian entries (GUE, GOE, Ginibre), the permanent follows an $\alpha$-stable distribution with stability index $\alpha\approx 1.0$$1.4$, well below the Gaussian value $\alpha=2$. We test Aaronson’s conjecture that $|\operatorname{perm}(X)|^{2}$ is asymptotically lognormal for Gaussian $X$: it is plausible for the complex Ginibre and GOE ensembles, but appears to fail for GUE and real Ginibre, where the $\alpha$-stable tails prevent convergence. Anti-concentration, however, holds for all Gaussian ensembles and is more robust than for Haar unitaries.
@@
for odd $n$ and zero for even $n$. We also study the geodesic forom the identity to the DFT matrix.
@@ 1. Introduction
  4. (4)
    Lognormal conjecture. We test Aaronson’s conjecture that $|\operatorname{perm}(X)|^{2}$ is asymptotically lognormal for Gaussian $X$. The conjecture is plausible for complex Ginibre and GOE, but appears to fail for GUE and real Ginibre. Anti-concentration holds—and is stronger than for Haar unitaries—across all Gaussian ensembles.
@@ 6.2. Failure of the CLT: $\alpha$-stable distributions
7. Aaronson’s lognormal conjecture
Aaronson and Arkhipov [1] proved that for the complex Ginibre ensemble ($X$ with i.i.d. $\mathcal{CN}(0,1)$ entries), $|\det(X)|^{2}$ converges to a lognormal random variable, and conjectured the same for $|\operatorname{perm}(X)|^{2}$ based on the visual similarity of their densities. We test this conjecture numerically for all four Gaussian ensembles.
7.1. Setup
For each ensemble (complex Ginibre, real Ginibre, GUE, GOE), we generate 50,00050{,}000 matrices at each $n$ and compute $\log|\cdot|$. Determinants are computed via LU decomposition for $n\in\{7,9,\ldots,101\}$; permanents via GPU-accelerated Ryser for $n\leq 17$. We test normality of $\log|\cdot|$ via Kolmogorov–Smirnov, Shapiro–Wilk, Anderson–Darling, and D’Agostino–Pearson tests, plus skewness and excess kurtosis.
7.2. Results
Table 4 compares $\log|\operatorname{perm}|$ and $\log|\det|$.
Table 4. Normality of $\log|\operatorname{perm}|$ vs. $\log|\det|$ (50,00050{,}000 samples).
$\log|\det|$ $\log|\operatorname{perm}|$
Ensemble $n$ Skew Kurt AD Skew Kurt AD
Ginibre-C 7 0.48-0.48 $0.61$ $111$ 0.48-0.48 $0.68$ $109$
13 0.38-0.38 $0.39$ $73$ 0.39-0.39 $0.65$ $68$
17 0.34-0.34 $0.34$ $55$ 0.39-0.39 $0.63$ $67$
GOE 7 0.63-0.63 $0.79$ $195$ 0.58-0.58 $0.87$ $154$
13 0.50-0.50 $0.60$ $126$ 0.47-0.47 $0.84$ $100$
17 0.44-0.44 $0.46$ $97$ 0.46-0.46 $0.93$ $87$
Ginibre-R 7 0.74-0.74 $1.34$ $249$ 0.77-0.77 $1.37$ $271$
13 0.66-0.66 $1.10$ $189$ 0.71-0.71 $1.37$ $217$
17 0.58-0.58 $0.89$ $147$ 0.77-0.77 $1.86$ $227$
GUE 7 0.76-0.76 $1.21$ $280$ 0.80-0.80 $1.56$ $285$
13 0.63-0.63 $1.05$ $171$ 0.69-0.69 $1.42$ $210$
17 0.57-0.57 $0.87$ $148$ 0.73-0.73 $1.86$ $211$
Figure 9 presents Q–Q plots comparing $\log|\det|$ and $\log|\operatorname{perm}|$ against normal quantiles for the complex Ginibre ensemble, and Figure 10 summarizes the skewness and kurtosis trajectories across all ensembles.
Figure 9. Q–Q plots of $\log|\det|$ (top, blue) and $\log|\operatorname{perm}|$ (bottom, red) against normal quantiles for the complex Ginibre ensemble. The determinant straightens visibly as $n$ grows from $7$ to $101$; the permanent improves more slowly over its accessible range $n\leq 17$.
Figure 10. Skewness and excess kurtosis of $\log|x|$ for determinants (top) and permanents (bottom) across all four Gaussian ensembles. Normal corresponds to $(0,0)$. The determinant converges toward $(0,0)$ for all ensembles; the permanent converges for GOE and Ginibre-C but not for GUE and Ginibre-R.
A striking dichotomy emerges:
  • Ginibre-C and GOE: The permanent tracks the determinant closely; both improve with $n$. Aaronson’s lognormal conjecture is plausible for these ensembles.
  • GUE and Ginibre-R: The permanent diverges from the determinant. While $\log|\det|$ steadily improves, $\log|\operatorname{perm}|$ shows flat or worsening skewness ($\approx{-}0.7$ to 0.8{-}0.8) and kurtosis ($\approx 1.4$ to $1.9$). A lognormal conjecture for these ensembles appears to be false.
7.3. Connection to $\alpha$-stable distributions
The divergent behavior of GUE/Ginibre-R permanents has a natural explanation via the $\alpha$-stable distributions of Section 6. For $\alpha<2$, the power-law tails $P(|X|>x)\sim x^{-\alpha}$ produce a heavy left tail in $\log|X|$ (many near-zero values), manifesting as persistent negative skewness that does not diminish with $n$. The hierarchy is: $\alpha\approx 1.4$ (Ginibre-C, lightest tails, lognormal most plausible) >> $\alpha\approx 1.2$ (GUE, Ginibre-R, lognormal doubtful) >> $\alpha\approx 1.0$ (GOE, Cauchy-like, yet $\log|\operatorname{perm}|$ still improves slowly). The determinant converges to lognormal for all ensembles because $\log|\det(X)|^{2}=\sum_{k}\log\chi^{2}_{2k}$ is a sum of independent random variables to which Berry–Esseen applies.
7.4. Anti-concentration for Gaussian ensembles
The PACC (see Section 8.3) asks whether $\Pr[|\operatorname{perm}|^{2}\geq\mathbb{E}[|\operatorname{perm}|^{2}]/\mathrm{poly}(n)]\geq 1/\mathrm{poly}(n)$. For Gaussian matrices, $\mathbb{E}[|\operatorname{perm}(X)|^{2}]=n!$ by independence of entries.
Observation 14.
Anti-concentration holds for all four Gaussian ensembles, and is in fact more robust than for Haar unitaries: at the threshold $\Pr[|\operatorname{perm}|^{2}\geq\mathbb{E}/n^{2}]$, the Gaussian ensembles show increasing probability with $n$ ($67\%$$98\%$ at $n=17$), while Haar unitaries decrease ($18\%$ at $n=17$).
The heavy $\alpha$-stable tails that destroy lognormality simultaneously strengthen anti-concentration: a symmetric $\alpha$-stable distribution with $\alpha>1$ has bounded density at the origin, preventing mass from piling up near zero. The failure of the CLT and the success of anti-concentration are two sides of the same coin (Figure 11).
Figure 11. Left: $\Pr[|\operatorname{perm}|^{2}\geq\mathbb{E}[|\operatorname{perm}|^{2}]/n^{2}]$ vs. matrix size $n$ for all ensembles. The Gaussian ensembles increase with $n$ while Haar unitaries decrease. Right: anti-concentration curve at $n=13$, showing $\Pr[|\operatorname{perm}|^{2}\geq\mathbb{E}/n^{k}]$ as a function of $k$.
@@ 8.4. Open problems
    Prove or disprove Aaronson’s lognormal conjecture for the complex Ginibre ensemble. For GUE and real Ginibre, our evidence suggests the conjecture is false.
  7. (7)